OpenStudy (anonymous):

Please Help !!; how should I Integrate ? double integral (xye^xydxdy [2,3]*[1,3]

OpenStudy (anonymous):

just notice \[\large \int\limits_{1}^{3} \int\limits_{2}^{3} xy e^{xy} dx dy = \int\limits_{1}^{3} y \ (\int\limits_{2}^{3} x e^{xy} dx) \ dy\]

OpenStudy (anonymous):

in the inner integral treat y like a constant

OpenStudy (cherylim23):

If you like or if it makes things easier, you can also swop the inside and outside terminals.